Calculate the change in energy (in units of kJ/mol) between the excited state and ground state for the transition that results in the emission of 285 nm light. (4 pts) A) 4.20 x 102 kJ/mol B) 6.20 x 102 kJ/mol C) 4.20 x 104 kJ/mol
Answer #1

Enho E is n is his & is the energy of radiation number of Photons Plancks constant = 6.626 X 16-34 J. See the frequency of rE= hho n = 1 mol = 6-023x10 23 6-0234 10 23 X 6.626 X 10 -345-secx 3x108 insert 985 10 9 m E = 420088.4 I E 4 20-08 11 12 E =